Use Gauss’ Law to calculate the electric field due to a long line of charge, with linear charge density 


1
Expert's answer
2021-04-28T20:19:41-0400

Answer

Let from to r distance of long line charge density "\lambda" where we want to calculate electric field.

So this situation consider a cylinderical shape as gaussian surface of radius r and length l

So using gauss law

"E\times2\pi rl=\frac{\lambda l}{\epsilon _0}"

Where "Q=\lambda l"

So electric field

"E=\frac{2k\lambda}{r}"